# A Design Framework for Distributed Quantum Computing Simulation

### v2

@@ 5 Conclusion and Future Work
There is a quite variegated choice of simulation tools for studying either quantum networks or quantum computers that may be used to support DQC research. Based on the proposed design framework, we are currently studying and developing modules that allow for bridging existing simulation tools, especially those that support seamless replacement of simulated hardware with real devices. For example, we plan to develop tools for orchestrating DQC simulations, with automated instantiation of simulation objects representing QPUs and quantum network components.
Acknowledgment
Michele Amoretti acknowledges financial support from the European Union – NextGenerationEU, PNRR MUR project PE0000023-NQSTI.

### v3

@@ 6 Conclusion and Future Work
There is a quite variegated choice of simulation tools for studying either quantum networks or quantum computers that may be used to support DQC research. Based on the proposed design framework, future work will concern the study and development of modules that allow for bridging existing simulation tools, with automated instantiation of simulation objects representing QPUs and quantum network components.
Acknowledgement
Davide Ferrari acknowledges financial support from the EU Flagship on Quantum Technologies through the project Quantum Internet Alliance (EU Horizon Europe, grant agreement no. 101102140). Michele Amoretti acknowledges financial support from the European Union – NextGenerationEU, PNRR MUR project PE0000023-NQSTI.